an arrow leaves a bow at 25m/s (a) what is the maximum range? (b) at what two angles above the the horizontal can it be sent to reach target 50m away?


1
Expert's answer
2022-02-15T16:14:18-0500

"l_{max}=\frac{v^2}g=62.5~m,"

"l=l_{max}\sin2\alpha,"

"\sin 2\alpha=\frac l{l_{max}}=\frac 45,"

"\alpha_1=26.6°,"

"\alpha_2=63.4°."